OpenStudy (anonymous):

Simplify x^6n - 1 Someone help me with this please the answer is (x^n+1)(x^n-1)(x^4n+x^2n +1)

OpenStudy (anonymous):

\[x^{6n}-1=(x^{2n})^3-(1)^3\] \[(x^{2n}-1)(x^{4n}-x^{2n}+1)=(x^n+1)(x^n-1)(x^{4n}-x^{2n}+1)\]

OpenStudy (anonymous):

hello...got it???

OpenStudy (anonymous):

Sort of one sec

OpenStudy (anonymous):

No I don't get it

OpenStudy (anonymous):

what is the problem??

OpenStudy (anonymous):

i first use the formula a^3-b^3=(a-b)(a^2+ab+b^2) and then (a^2-b^2)=(a+b)(a-b)

OpenStudy (anonymous):

thank you i will try to work it out right now